Are the following statements true or false? Give reasons for tour answers.

a) −2 isalimitpointoftheinterval ]−3,2].

b) The series (1/2) - (1/6) + (1/10) (−1/4) +.... is divergent.

c) The function, f (x) = sin2x is uniformly continuous in the interval [0,π].

d) Every continuous function is differentiable.

e) The function f defined on R by

f(x)= {0, if x is rational and 2, if x is irrational

Is integral element in the interval [2,3].


1
Expert's answer
2021-06-07T12:00:31-0400

a) true

Notation "(a,b]" means that "b" is included in the interval.


b) false

Since "\frac{|a_{n+1}|}{|a_n|}<1" , the series is convergent.


c) false

 The function f(x) is uniformly continuous in the interval "[0,2\pi]" if

"\forall\ \varepsilon\ \exists\ \delta>0\ \forall\ x_1,x_1\isin[0,2\pi]\ (|x_1-x_2|<\delta)\implies (f(x_1)-f(x_2)<\varepsilon)"

"Let\ \varepsilon=1" and "sin^2x_1=0,sin^2x_2=\pm 1"

Then:

"x_1^2=\pi n,x_2=\pi n+\pi/2"


"|x_1-x_2=\sqrt{\pi n+\pi/2}-\sqrt{\pi n}=\frac{\pi/2}{\sqrt{\pi n+\pi/2}+\sqrt{\pi n}}<\frac{2}{2\sqrt{\pi n}}<\frac{1}{\sqrt{n}}"

If "n>1/\delta^2" then "|x_1-x_2|<\delta" , but"|f(x_1)-f(x_2)|=1"

So, the given function is not uniformly continuous in the interval [0,π].


d) false

For example, continuous function "f(x)=|x|" is not differentiable at x=0


e) false

An element b of a commutative ring B is said to be integral over A, a subring of B, if there are n ≥ 1 and aj in A such that

"b^n+a_{n-1}b^{n-1}+...+a_1b+a_0=0"

For the given function:

 if x is irrational, then there are no coefficients "x_0,x_1,...,x_n" , such that:

"2^n+2^{n-1}\cdot x_{n-1}+...+2x_1+x_0=0"